Map apling losming The half-lives of different radioisotopes are given in the table below. How long would i take (in minutes) for the amount of lead-196 to decrease from,20.0 mg to 2.50 mg? w. Howlong would ake (in minutes, Radioisotope Half-life (min) argon 44 ead-196 potassium-44 indium-117 12 37 43 Number 0 min Tocis
Answer #1


N = No*(1/2)^(t/T1/2)

T1/2 = half life period = 37 min

No = iniital amount

t = time required


N = remaining substance

2.5 = 20*(1/2)^(t/37)

t = 111 min